ln(x + 4) + ln(x) = ln(x + 4

sovle for equation for x
Answer #1

* Given equation is In(x+4) fIn(a) = ln (x+4). we have ln m + lnn = ln (mn). In x(x64) - In(2+4). o x(x+4) = x44 = a + 4X = X